There is a question close to this on the board but it is slightly different. Can you please help me.
Thank you.
Solve using the quadratic formula:
3x^2-2x=15x-10

Answer by stanbon(75887)

Solve using the quadratic formula:
3x^2-2x=15x-10
3x^2-17x+10 = 0
3x^2-15x-2x+10 = 0
3x(x-5)-2(x-5)=0
(x-5)(3x-2)=0
x = 5 or x = 2/3
======================
Cheers,
Stan H.